OpenStudy (anonymous):

solve for s: 1/r=1/s+1/t

OpenStudy (jdoe0001):

use the LCD again :)

OpenStudy (jdoe0001):

$$ \cfrac{1}{r}=\cfrac{1}{s}+\cfrac{1}{t} \implies \cfrac{1}{r}-\cfrac{1}{t}=\cfrac{1}{s}\\ \cfrac{1}{r}-\cfrac{1}{t} \color{blue}{\ add\ this}\\ \text{when done multiply both sides by }\color{blue}{s} $$

OpenStudy (jdoe0001):

hmm... one sec

OpenStudy (jdoe0001):

well, you'd need to multiply for more than just "s"

OpenStudy (jdoe0001):

$$ \cfrac{1}{r}=\cfrac{1}{s}+\cfrac{1}{t} \implies \cfrac{1}{r}-\cfrac{1}{t}=\cfrac{1}{s}\\ \cfrac{1}{r}-\cfrac{1}{t} \color{blue}{\ add\ this}\\ \text{when done multiply both sides by }\\ \cfrac{\boxed{\alpha}}{\boxed{\theta}}=\cfrac{1}{s} \color{blue}{\times \cfrac{\boxed{\theta}}{\boxed{\alpha}}\times \cfrac{s}{1}} $$

OpenStudy (anonymous):

theansweri got was \[s=\frac{ 1 }{ (1/t)-(1/r)}\]

OpenStudy (jdoe0001):

lemme check

OpenStudy (jdoe0001):

in rational form, it'd be $$ \large s=\cfrac{1}{\frac{1}{r}-\frac{1}{t}} $$

OpenStudy (jdoe0001):

but I assume you're still expected to further rationalize the bottom fraction

OpenStudy (anonymous):

no idont think i do

OpenStudy (jdoe0001):

ok :)